Session: Jun 2010 (session 2)
Wavelength: 5 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (Orfei at Tue Jun 01 15:09:45 2010) Minor failures:
VC02 had some locking problem.

Yebes (Paco Colomer at Tue Jun 01 15:33:37 2010) Failure: ANT
The antenna cannot be moved. This experiment has not been observed.

Torun (K. Borkowski at Wed Jun 02 06:46:48 2010) Success:
No known problems apart from bit statistics which were not as one would wish (however we cannot do much to rectify this situation).

Effelsberg (Uwe Bach at Wed Jun 02 09:10:50 2010) Success:
First 5 min were lost because of wrong IF connected.

Onsala (Michael Lindqvist at Thu Jun 03 07:01:50 2010) Minor failures:
Mark 5 problems first 10 minutes. Otherwise no known problems.

Westerbork (Geert Kuper at Fri Jun 04 08:17:29 2010) Success:
Only RT7 in TADUmax

Shanghai (Bo Xia at Wed Jun 09 07:32:51 2010) Success: